OpenStudy (anonymous):

A hair dryer draws 14.5A when plugged into a 120-V line. Assume direct current. A) What is its resistance? for this i got 8.3 (ohm) but i dont know what to do with the second part of the question B) How much charge passes through it in 15min ?

OpenStudy (anonymous):

for Part a i did R = V/ A

hartnn (hartnn):

Charge = current * time part a is correct :)

OpenStudy (anonymous):

I dont understand which formula to use for part b.

hartnn (hartnn):

i gave you the formula only Charge = current * time current = 14.5 time = 15 min

hartnn (hartnn):

if needed convert time into seconds time = 15 min = 15*60 sec = ....sec

OpenStudy (anonymous):

oh that is all, I didnt even find that in the book thank you very much

hartnn (hartnn):

your most welcome ^_^

OpenStudy (anonymous):

I got an answer of 13050

hartnn (hartnn):

charge = 13050 Ampere-sec don't forget the units :)

OpenStudy (anonymous):

so it would be A/s ..your the best thank you!

hartnn (hartnn):

not A/s its A-s A (ampere ) and s (seconds) are multiplied, not divided

hartnn (hartnn):

because charge is current(in amperes) multiplied by time(in seconds)

OpenStudy (anonymous):

A*s right, lemme try it aha thank you!

OpenStudy (anonymous):

it worked thank you!

hartnn (hartnn):

:D